OpenStudy (anonymous):

PLEASE HELP!! Solve for extraneous solution sqrt. sqrt. x+25 = sqrt. x+5

OpenStudy (anonymous):

u need help solving it

OpenStudy (anonymous):

yes please! D: this is the last equation

OpenStudy (anonymous):

use this web http://www.purplemath.com/modules/solverad2.htm

OpenStudy (anonymous):

http://www.sosmath.com/CBB/viewtopic.php?t=2745

OpenStudy (anonymous):

kay, thank you ^^ I'll try.. It confuses me because theres double square roots, like this \[\sqrt{\sqrt{x+25}}=\sqrt{x+5}\]

OpenStudy (anonymous):

ummm i want u too really underdstanf it did it help

OpenStudy (anonymous):

tell me its k

OpenStudy (anonymous):

okay that last link helps! Thank you!!!

OpenStudy (anonymous):

ur welcome i am only in 5th grade gifted class

OpenStudy (anonymous):

and i medal fan u